# Incentive Mechanism Design for Privacy-Preserving Decentralized Blockchain Relayers

### v1

@@ Abstract
Keywords Blockchain $\cdot$ Privacy $\cdot$ Relayers $\cdot$ Game-Theory $\cdot$ Evolutionary Games $\cdot$ Volunteer’s Dilemma
@@ 1 Introduction
  • We establish the existence, uniqueness, and evolutionary stability of the symmetric mixed-strategy Nash equilibrium of the game.
@@ 2.1 Privacy in Public Blockchains
Privacy in public blockchains has been addressed through several complementary approaches that operate at different layers. Cryptographic mechanisms, such as zero-knowledge proofs and threshold credential systems, provide strong unlinkability guarantees at the application or data layer but introduce significant computational overhead citezerocash2014,nymCredentials. Network-layer anonymity mechanisms aim to obscure transaction origin during propagation; protocols such as Dandelion++ and its subsequent analyses rely on controlled gossiping and probabilistic routing to reduce linkability [5, 7]. More recent designs combine cryptographic techniques with layered routing, as in mixnet-based architectures such as Loopix, where batching and cover traffic further strengthen sender anonymity [22]. While each of these approaches mitigates specific privacy risks, none is sufficient on its own, and practical systems often combine techniques across layers to address complementary threats.
@@ 4.2 Equilibrium Analysis for $\mathcal{G}_{N}$
Consider the function $g(q_{u})=v(\mathrm{U},q_{u})-v(\mathrm{NU},q_{u}),$ which represents the expected utility gain from uploading when all other relayers upload with probability $q_{u}$. When $q_{u}=0$, no other relayer uploads, and a unilateral upload strictly improves the outcome, so $g(0)>0$. When $q_{u}=1$, all relayers upload, and uploading yields only its cost, implying $g(1)<0$. Since $g$ is continuous, there exists at least one root in $(0,1)$. Substituting the explicit utilities yields the polynomial $h(q_{u})$ given in Theorem 1. An analysis of its derivative shows that $h^{\prime}(q_{u})$ changes sign exactly once on $(0,1)$, so $h$ attains a single local maximum and crosses zero at a unique point. This point corresponds to the unique symmetric mixed-strategy Nash equilibrium $q_{u}^{*}$. The proof is provided in Appendix A.
@@ 5.1 Evaluation Parameters and Metrics
To evaluate the proposed model, we vary the main system parameters and measure their effect on the upload probability and overall system performance at equilibrium. Throughout all simulations, the benefit is fixed at $b=100$, while the other parameters are varied as a ratio of the benefit. These parameters include: the cost of the first upload $c_{f}$, the cost of later uploads $c_{l}$ the penalty for a no-upload event $p$, and the number of relayers $N$. Since in practice, successful execution of method calls incur higher cost than reverted ones, $c_{l}$ is chosen as a unit of cost. The ratio between $c_{f}$ and $c_{l}$ varies depending the actual code executed. The numerical values are normalized rather than literal gas measurements; only the ratios between $b,c_{f},c_{l}$, and $p$ influence equilibrium behavior.
@@ 5.2 Numerical Results
5.2.1 Estimation of the cost-reward ratio
In order to parametrize the numerical evaluation, as b and p are system paramaters controlled by system designers, it is more important to estimate the ratio of $c_{l}$ and $c_{f}$. As stated before, $c_{f}$ captures the cost in terms of units of cost of a successful execution of the transaction of upload. While $c_{l}$ is the cost for a reverted transaction. In order to estimate the magniture between the two. We implement three basic smart contracts in solidity and call a successful execution and a reverted one. in this case, the
@@ A.1 Proof of Theorem 1
Non-existence of symmetric pure Nash equilibria. When all players choose not to upload, then the best response for any individual player is to deviate and upload, changing their utility from p<0-p<0 to $r-c_{f}>0$. Similarly, if all players choose to upload, then the best response is to not upload, and thus receiving a reward of $b$, which is preferable to either $b-c_{f}$ or $b-c_{l}$. Therefore, in the symmetric case, no pure strategy equilibrium exists.
@@ A.2 Proof of Proposition 1
To begin, we note that $\mathcal{G}_{N}$ is a symmetric game. Indeed, all players share the same action space, and the payoff functions are invariant under any permutation of players’ indices and actions. Furthermore, since each player has exactly two strategies, it is known that any symmetric $N$-player game with a common two-strategy set is an exact potential game [11, Section 4.2.2].
@@
Moreover, since $\mathcal{G}_{N}$ is a potential game, the asymptotic stability of the Nash equilibrium can be established by invoking a standard result from evolutionary game theory: any isolated local maximizer of a potential function is asymptotically stable under any evolutionary dynamic that satisfies positive correlation and Nash stationarity [23, Theorem 8.2.1].
@@ A.4 Proof of Corollary 1
$W(a^{*})$ is the maximum reward that could be distributed to the players. Hence, every pure NE is a global maximizer of $W$. Interestingly, this also implies that every pure NE is a strong equilibrium, as any socially optimal nash equilibrium is a strong equilibrium [13, Theorem 33].

### v2

@@ Abstract
Keywords Blockchain $\cdot$ Privacy $\cdot$ Relayers $\cdot$ Game Theory $\cdot$ Evolutionary Games $\cdot$ Volunteer’s Dilemma
@@ 1 Introduction
  • We establish the existence, uniqueness, and evolutionary stability of the symmetric mixed-strategy Nash equilibrium (NE) of the game.
@@ 2.1 Privacy in Public Blockchains
Privacy in public blockchains has been addressed through several complementary approaches that operate at different layers. Cryptographic mechanisms, such as zero-knowledge proofs and threshold credential systems, provide strong unlinkability guarantees at the application or data layer but introduce significant computational overhead [1, 10]. Network-layer anonymity mechanisms aim to obscure transaction origin during propagation; protocols such as Dandelion++ and its subsequent analyses rely on controlled gossiping and probabilistic routing to reduce linkability [6, 8]. More recent designs combine cryptographic techniques with layered routing, as in mixnet-based architectures such as Loopix, where batching and cover traffic further strengthen sender anonymity [23]. While each of these approaches mitigates specific privacy risks, none is sufficient on its own, and practical systems often combine techniques across layers to address complementary threats.
@@ 4.2 Equilibrium Analysis for $\mathcal{G}_{N}$
Consider the function $g(q_{u})=v(\mathrm{U},q_{u})-v(\mathrm{NU},q_{u}),$ which represents the expected utility gain from uploading when all other relayers upload with probability $q_{u}$. When $q_{u}=0$, no other relayer uploads, and a unilateral upload strictly improves the outcome, so $g(0)>0$. When $q_{u}=1$, all relayers upload, and uploading yields only its cost, implying $g(1)<0$. Since $g$ is continuous, there exists at least one root in $(0,1)$. Substituting the explicit utilities yields the polynomial $h(q_{u})$ given in Theorem 1. An analysis of its derivative shows that $h^{\prime}(q_{u})$ changes sign exactly once on $(0,1)$, so $h$ attains a single local maximum and crosses zero at a unique point. This point corresponds to the unique symmetric mixed-strategy Nash equilibrium $q_{u}^{*}$. The proof is provided in Appendix A.1.
@@ 5.1 Evaluation Parameters and Metrics
To evaluate the proposed model, we vary the main system parameters and measure their effect on the upload probability and overall system performance at equilibrium. Throughout all simulations, the benefit is fixed at $b=100$, while the other parameters are varied as a ratio of the benefit. These parameters include: the cost of the first upload $c_{f}$, the cost of later uploads $c_{l}$ the penalty for a no-upload event $p$, and the number of relayers $N$. Since in practice, successful execution of method calls incurs higher cost than reverted ones, $c_{l}$ is chosen as a unit of cost. The ratio between $c_{f}$ and $c_{l}$ varies depending the actual code executed. The numerical values are normalized rather than literal gas measurements; only the ratios between $b,c_{f},c_{l}$, and $p$ influence equilibrium behavior.
@@ A.1 Proof of Theorem 1
Non-existence of symmetric pure Nash equilibria. When all players choose not to upload, then the best response for any individual player is to deviate and upload, changing their utility from p<0-p<0 to $b-c_{f}>0$. Similarly, if all players choose to upload, then the best response is to not upload, and thus receiving a reward of $b$, which is preferable to either $b-c_{f}$ or $b-c_{l}$. Therefore, in the symmetric case, no pure strategy equilibrium exists.
@@ A.2 Proof of Proposition 1
To begin, we note that $\mathcal{G}_{N}$ is a symmetric game. Indeed, all players share the same action space, and the payoff functions are invariant under any permutation of players’ indices and actions. Furthermore, since each player has exactly two strategies, it is known that any symmetric $N$-player game with a common two-strategy set is an exact potential game [12, Section 4.2.2].
@@
Moreover, since $\mathcal{G}_{N}$ is a potential game, the asymptotic stability of the Nash equilibrium can be established by invoking a standard result from evolutionary game theory: any isolated local maximizer of a potential function is asymptotically stable under any evolutionary dynamic that satisfies positive correlation and Nash stationarity [24, Theorem 8.2.1].
@@ A.4 Proof of Corollary 1
$W(a^{*})$ is the maximum reward that could be distributed to the players. Hence, every pure NE is a global maximizer of $W$. Interestingly, this also implies that every pure NE is a strong equilibrium, as any socially optimal Nash equilibrium is a strong equilibrium [14, Theorem 33].